# Horizon Brightened Acceleration Radiation from Massive Vector Fields

### v1

@@ Abstract
In this paper, we develop a quantum-optical treatment of acceleration radiation for atoms freely falling into a Schwarzschild black hole when the ambient field is a massive spin-1 (Proca) field. Building on the HBAR framework of Scully and collaborators, we analyze two detector realizations: a charged-monopole current coupling and a physical electric-dipole coupling, both within a cavity that isolates a single outgoing Schwarzschild mode prepared in the Boulware state. Using a near-horizon stationary-phase analysis, we show that the thermal detailed-balance factor governing excitation versus absorption is universal and depends only on the near-horizon Rindler coordinate transformation. At the same time, the absolute spectra acquire distinctive Proca signatures: a hard mass threshold, polarization-dependent prefactors, and axial/polar greybody transmissions. Promoting single-pass probabilities to escaping rates yields a master equation whose steady state is geometric and whose entropy flux obeys an HBAR-style area-entropy relation identical in form to the scalar case, with all vector-field specifics entering through the radiative area change. Our results provide a controlled pathway to probe longitudinal versus transverse responses, mass thresholds, and polarization-resolved greybody effects in acceleration radiation, and set the stage for extensions to rotating backgrounds, alternative exterior states, and detector-engineering strategies.
@@ I Introduction
Black holes provide a unique meeting point between gravitation, quantum field theory, and statistical physics. Seminal work showed that quantum fields in curved spacetime endow black holes with a thermal character, leading to particle emission and an associated thermodynamics governed by horizon geometry Bekenstein:1973ur; Hawking:1974rv; Hawking:1975vcx. The modern picture originates with Hawking’s prediction of black‐hole radiation and the area-entropy paradigm that followed, which together frame today’s discussions of information, equilibration, and backreaction in strong gravity.
@@
Beyond spontaneous emission at future null infinity, accelerated motion in vacuum can itself stimulate radiation. In flat spacetime this appears as the Unruh effect, for which accelerated detectors register a thermal spectrum set by their proper acceleration; comprehensive accounts have clarified both foundational aspects and operational detector models Unruh:1976db. In black‐hole spacetimes, the same near‐horizon mapping that underlies Hawking’s result organizes the response of localized probes and the structure of physically relevant vacuum states Boulware:1974dm
@@
A particularly incisive formulation of this physics, which is directly relevant for quantum‐optical probes, is the horizon brightened acceleration radiation (HBAR) program introduced by Scully and collaborators Scully:2017utk. In that setting, atoms prepared in their ground state fall through a cavity toward a Schwarzschild horizon while the exterior field is in the Boulware vacuum; the atoms emit acceleration radiation via a counter-rotating channel, and a master‐equation treatment yields an entropy flux that closely parallels black‐hole thermodynamic relations. Subsequent work developed this quantum‐optics viewpoint and emphasized its complementarity to conventional Hawking derivations Azizi:2021qcu; Azizi:2021yto. Since then, several studies have emerged to use accelerated radiation to probe certain geometric signatures from different black hole models Svidzinsky:2018jkp; Ben-Benjamin:2019opz; Azizi:2020gff; Azizi:2021qcu; Azizi:2021yto; Camblong:2020pme; Sen:2022cdx; Sen:2022tru; Das:2023rwg; Das:2025rzz; Jana:2024fhx; Jana:2025hfl; Bukhari:2022wyx; Bukhari:2023yuy; Ovgun:2025isv; Pantig:2025okn; Pantig:2025igg, and even on time-dependent black hole metrics Dalui:2020qpt; Kaplanek:2021sbo; Shallue:2025zto; Clarke:2024lwi; Carullo:2025oms.
@@
Motivated by both conceptual and phenomenological considerations, we revisit this scenario for a massive spin-1 field governed by the Proca equations. Massive vector fields introduce qualitatively new ingredients: a longitudinal polarization with distinct near-horizon couplings; a hard spectral threshold set by the rest mass that reshapes low-frequency response; and polarization-dependent propagation through curvature-induced barriers that control the greybody transmission to infinity. These features are well studied in black-hole perturbation theory, where the Proca field separates into axial and polar sectors with one and two effective channels, respectively, and displays characteristic quasinormal and quasibound spectra Konoplya:2005hr; Rosa:2011my; Fernandes:2021qvr. They also connect to broader interest in hidden photons and dark sectors, for which black-hole environments provide sensitive laboratories.
@@
From the propagation side, the Schwarzschild problem for Proca fields has been analyzed in detail, including decoupled master equations, axial/polar structure, and numerical spectra; recent work has begun to systematize greybody factors for massive vectors, underscoring the importance of transmission coefficients in any observational or analogue setup Rosa:2011my. These developments make the Proca generalization of the HBAR framework both timely and natural: longitudinal versus transverse responses, mass thresholds, and polarization-specific greybody factors jointly provide clean spectral signatures unavailable to scalar fields. More broadly, interest in light vectors around black holes, especially in the context of ultralight hidden photons, has emphasized superradiant dynamics and observational signatures, underscoring the phenomenological value of polarization- and mass-threshold effects in spin-1 spectra Brito:2015oca; Baryakhtar:2017ngi. Finally, the role of cavities and localized detectors in curved backgrounds has been analyzed in detector-theory language, offering complementary perspectives on how boundary conditions and motion shape response, which are concepts directly mirrored by the cavity protocol adopted here Ahmadzadegan:2013iua.
@@
We adopt Scully et al.’s Scully:2017utk kinematic and operational scaffold (radial free fall into a Schwarzschild black hole, Boulware exterior state, and a mode-selecting cavity) to isolate an outgoing channel and evaluate the detector response. Our analysis keeps the detector modeling explicit: we treat (i) a charged-monopole current coupling and (ii) a physical electric-dipole coupling to the local field strength, thereby bracketing the range of realistic atom-field interactions in this geometry. Throughout, we emphasize what is universal (the near-horizon phase structure and the resulting detailed balance) versus what is model-dependent (polarization projectors, mass thresholds, and greybody transmission).
@@
We organize the paper as follows: Section II fixes the physical setting and kinematics of radial infall, the exterior vacuum, and frequency labels. Section III formulates the Proca dynamics on Schwarzschild, including separation into axial/polar sectors and mode normalization. Section IV specifies the two detector models and derives the corresponding worldline amplitudes. Section V evaluates the excitation probability using a near-horizon stationary-phase analysis, isolating the universal thermal kernel and identifying the Proca-specific prefactors without invoking explicit formulas here. Section VI incorporates the cavity protocol and greybody transmission to promote probabilities to emission/absorption rates and to a master equation for the mode occupations. Section VII develops the entropy-flux analysis (HBAR) for massive spin-1 quanta and relates it to the black-hole area change. Section VIII assembles consistency checks and theoretical subtleties, including control of the large-gap expansion, the role of the Proca constraint in place of gauge freedom, and the connection to Unruh/DeWitt detector logic. We close in Section IX with a brief summary and outlook.
@@ II Preliminaries and Conventions
We consider two-level atoms falling radially from rest at infinity into a nonrotating Schwarzschild black hole of mass $M$. The exterior quantum field is prepared in the Boulware vacuum. A mode-selecting cavity isolates counter-propagating outgoing modes so that the atom-field interaction reduces to a single relevant frequency label $\nu$ for distant observers. Kinematics is encoded by the Schwarzschild line element and the tortoise coordinate $r_{*}$, and the atomic worldline is specified by $t(\tau)$ and $r(\tau)$ for proper time $\tau$. The near-horizon redshift makes the WKB phase of outgoing modes depend on the combination $t-r_{*}$, which will be the driver of the thermal detailed-balance factor later on.
@@ II.1 Geometry, trajectory, and coordinates
We work outside the Schwarzschild radius $r_{g}\equiv 2GM/c^{2}$ with line element Schwarzschild:1916uq; Chandrasekhar:1985kt
$ds^{2}=f(r)\,c^{2}dt^{2}-\frac{dr^{2}}{f(r)}-r^{2}(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}),\qquad f(r)=1-\frac{r_{g}}{r}.$ (1)
For wave propagation, it is convenient to introduce the Regge-Wheeler tortoise coordinate Regge:1957td; Fiziev:2007es
$r_{*}(r)=r+r_{g}\ln\!\left(\frac{r}{r_{g}}-1\right),$ (2)
so that outgoing high-frequency modes behave locally as $e^{-i\nu(t-r_{*}/c)}$ near the horizon (the precise field content will be specified later). In a small neighborhood of the horizon, one may use the Rindler approximation obtained by setting $r=r_{g}+z^{2}/(4r_{g})$ with $0<z\ll r_{g}$, which yields Birrell:1982ix
$ds^{2}\simeq\frac{z^{2}}{4r_{g}^{2}}\,c^{2}dt^{2}-dz^{2}-r_{g}^{2}d\Omega_{2}^{2},$ (3)
@@
Following Ref. Scully:2017utk, we adopt the dimensionless variables
$r\rightarrow r_{g}\,r,\qquad t\rightarrow(r_{g}/c)\,t,\qquad\omega\rightarrow(c/r_{g})\,\omega,\qquad\nu\rightarrow(c/r_{g})\,\nu,$ (4)
@@
$\frac{dr}{d\tau}=-\frac{1}{\sqrt{r}},\qquad\frac{dt}{d\tau}=\frac{r}{r-1}\;,$ (5)
where $t$ is the Schwarzschild time and $\tau$ the atom’s proper time (both dimensionless in the convention of (4)). Integrating (5), we obtain the parametric trajectory
$\tau(r)=-\frac{2}{3}\,r^{3/2}+{\rm const},$ (6)
$t(r)=-\frac{2}{3}\,r^{3/2}-2\sqrt{r}-\ln\!\left(\frac{\sqrt{r}-1}{\sqrt{r}+1}\right)+{\rm const},$ (7)
and the corresponding tortoise coordinate Regge:1957td; Fiziev:2007es
$r_{*}(r)=r+\ln(r-1).$ (8)
Equations (5)-(8) completely specify $t(\tau)$, $r(\tau)$, and $r_{*}(\tau)$ for the ensuing stationary-phase analysis of the atom-field matrix elements. These expressions coincide with the kinematic relations used in Appendices A and B of Scully et al. Scully:2017utk, to which we will refer for consistency of conventions.
We remark that the near-horizon Rindler form (3) rationalizes the universality of the phase $t-r_{*}$ and will underpin the Planckian factor arising later from the $\tau$-integral along the worldline. 2) The choice of additive constants in (6)-(7) fixes the zero of $t$ and $\tau$ and is immaterial for transition probabilities, which depend only on phase differences.
@@ II.2 Field state and cavity
We take the exterior quantum field to be in the Boulware vacuum, i.e. the vacuum defined by positive Killing frequency with respect to $\partial_{t}$ in the static exterior. Operationally, we place a one-dimensional, axially symmetric cavity along the radial direction to isolate a single, outgoing mode counterpropagating with the infalling atom cloud (the same role played by the cavity/mode selector in Scully et al Scully:2017utk). The cavity suppresses contamination by incoming flux and renders the coarse-grained master-equation description well-posed; leakage can be neglected provided the photon loss rate is small compared to the atom-induced absorption rate.
@@
$a_{\nu,\lambda}\,|0_{B}\rangle=0,\qquad\nu>0,$ (11)
where $\lambda$ labels polarization (for the Proca field) or the single scalar polarization Boulware:1974dm; Candelas:1980zt. The near-horizon WKB form of each outgoing mode depends on the universal phase $t-r_{*}$; hence, when the atom falls through the cavity, the dominant contribution to the transition amplitude is governed by this phase, while the cavity merely enforces mode selection and provides the coarse graining needed for a laser-style master equation and the ensuing HBAR entropy analysis.

### v2

@@ Abstract
In this paper, we develop a quantum-optical treatment of acceleration radiation for atoms freely falling into a Schwarzschild black hole when the ambient field is a massive spin-1 (Proca) field. Building on the HBAR framework of Scully and collaborators, we analyze two detector realizations: a charged-monopole current coupling and a physical electric-dipole coupling, both within a cavity that isolates a single outgoing Schwarzschild mode prepared in the Boulware state. Using a near-horizon stationary-phase analysis, we show that the thermal detailed-balance factor governing excitation versus absorption is universal and depends only on the near-horizon Rindler coordinate transformation. At the same time, the absolute spectra acquire distinctive Proca signatures: a hard mass threshold, polarization-dependent prefactors, and axial/polar graybody transmissions. Promoting single-pass probabilities to escaping rates yields a master equation whose steady state is geometric and whose entropy flux obeys an horizon brightened acceleration radiation-style area-entropy relation identical in form to the scalar case, with all vector-field specifics entering through the radiative area change. Our results provide a controlled pathway to probe longitudinal versus transverse responses, mass thresholds, and the role of polarization-resolved graybody transmission in acceleration radiation. More precisely, we derive the universal near-horizon kernel and show how the Proca transmission data enter the escaping probabilities, rates, and entropy flux; a dedicated numerical computation of the axial/polar graybody profiles is left for future work. This sets the stage for extensions to rotating backgrounds, alternative exterior states, and detector-engineering strategies.
@@ I Introduction
Black holes provide a unique meeting point between gravitation, quantum field theory, and statistical physics. Seminal work showed that quantum fields in curved spacetime endow black holes with a thermal character, leading to particle emission and an associated thermodynamics governed by horizon geometry [7, 28, 29]. The modern picture originates with Hawking’s prediction of black‐hole radiation and the area-entropy paradigm that followed, which together frame today’s discussions of information, equilibration, and backreaction in strong gravity.
@@
Beyond spontaneous emission at future null infinity, accelerated motion in vacuum can itself stimulate radiation. In flat spacetime this appears as the Unruh effect, for which accelerated detectors register a thermal spectrum set by their proper acceleration; comprehensive accounts have clarified both foundational aspects and operational detector models [51]. In black‐hole spacetimes, the same near‐horizon mapping that underlies Hawking’s result organizes the response of localized probes and the structure of physically relevant vacuum states [10]
@@
A particularly incisive formulation of this physics, which is directly relevant for quantum‐optical probes, is the horizon brightened acceleration radiation (HBAR) program introduced by Scully and collaborators [44]. In that setting, atoms prepared in their ground state fall through a cavity toward a Schwarzschild horizon while the exterior field is in the Boulware vacuum; the atoms emit acceleration radiation via a counterrotating channel, and a master‐equation treatment yields an entropy flux that closely parallels black‐hole thermodynamic relations. Subsequent work developed this quantum‐optics viewpoint and emphasized its complementarity to conventional Hawking derivations [3, 4]. Since then, several studies have emerged to use accelerated radiation to probe certain geometric signatures from different black hole models [48, 8, 2, 3, 4, 15, 46, 45, 24, 23, 30, 31, 13, 14, 37, 39, 38], and even on time-dependent black hole metrics [22, 32, 47, 20, 18].
@@
Motivated by both conceptual and phenomenological considerations, we revisit this scenario for a massive spin-1 field governed by the Proca equations. Massive vector fields introduce qualitatively new ingredients: a longitudinal polarization with distinct near-horizon couplings; a hard spectral threshold set by the rest mass that reshapes low-frequency response; and polarization-dependent propagation through curvature-induced barriers that control the greybody transmission to infinity. These features are well studied in black hole perturbation theory, where the Proca field separates into axial and polar sectors with one and two effective channels, respectively, and displays characteristic quasinormal and quasibound spectra [33, 41, 25]. They also connect to broader interest in hidden photons and dark sectors, for which black hole environments provide sensitive laboratories.
@@
From the propagation side, the Schwarzschild problem for Proca fields has been analyzed in detail, including decoupled master equations, axial/polar structure, and numerical spectra; recent work has begun to systematize greybody factors for massive vectors, underscoring the importance of transmission coefficients in any observational or analogue setup [41]. These developments make the Proca generalization of the HBAR framework both timely and natural. The physical interest is not merely that one obtains a spin-1 version of the scalar calculation, but that massive vectors introduce qualitatively new observables: a longitudinal polarization absent in scalar and Maxwell cases, a hard spectral threshold at $\nu=m_{V}$, and polarization-resolved propagation through axial/polar Schwarzschild barriers. Within a cavity-selected HBAR protocol, these features imply a distinctive gap-plus-tail spectrum, a channel-dependent turn-on near threshold, and detector-dependent sensitivity to longitudinal versus transverse response. These are semirealistic signatures in the sense relevant to detector theory and analogue implementations: they specify which observable features would distinguish a massive vector field from scalar or massless spin-1 radiation in any mode-resolved measurement. More broadly, because light vectors and hidden photons are actively studied in black hole environments, the Proca case provides a useful analytic template for separating universal near-horizon thermality from field-specific spin-1 spectral structure. More broadly, interest in light vectors around black holes, especially in the context of ultralight hidden photons, has emphasized superradiant dynamics and observational signatures, underscoring the phenomenological value of polarization- and mass-threshold effects in spin-1 spectra [12, 6]. Finally, the role of cavities and localized detectors in curved backgrounds has been analyzed in detector-theory language, offering complementary perspectives on how boundary conditions and motion shape response, which are concepts directly mirrored by the cavity protocol adopted here [1].
@@
We adopt Scully et al.’s [44] kinematic and operational scaffold (radial free fall into a Schwarzschild black hole, Boulware exterior state, and a mode-selecting cavity) to isolate an outgoing channel and evaluate the detector response. This choice should be understood operationally rather than as a claim about the globally physical late-time state of an evaporating black hole. In the present HBAR protocol, the role of the Boulware vacuum is to provide a zero-occupation reference state for the cavity-filtered outgoing mode, so that the atom’s excitation isolates the acceleration-radiation channel without contamination from a preexisting thermal bath. Our analysis keeps the detector modeling explicit: we treat (i) a charged-monopole current coupling and (ii) a physical electric-dipole coupling to the local field strength, thereby bracketing the range of realistic atom-field interactions in this geometry. Throughout, we emphasize what is universal (the near-horizon phase structure and the resulting detailed balance) versus what is model-dependent (polarization projectors, mass thresholds, and greybody transmission). In the present work, the greybody sector is incorporated at the level of general Schwarzschild scattering data: we identify the transmission coefficients that enter the cavity-selected HBAR rates, but we do not attempt a separate numerical computation of the axial/polar Proca greybody profiles. Thus, our main result is a factorized analytic framework in which the local near-horizon kernel is derived explicitly, while the global propagation sector enters through transmission inputs.
We stress, however, that the Boulware state is an idealized baseline for the cavity-selected problem. For an evaporating black hole, the more physical exterior state is the Unruh vacuum, whereas the Hartle-Hawking state describes thermal equilibrium. In those alternatives, the same near-horizon worldline kernel derived below remains the local building block, but the selected mode is no longer initially empty, so the effective emission/absorption rates acquire additional state-dependent Bose factors. The present paper therefore isolates the vacuum-seeded HBAR channel; a systematic incorporation of Unruh/Hartle-Hawking occupation effects is left for future work.
@@
We organize the paper as follows: Sec. II fixes the physical setting and kinematics of radial infall, the exterior vacuum, and frequency labels. Sec. III formulates the Proca dynamics on Schwarzschild, including separation into axial/polar sectors and mode normalization. Sec. IV specifies the two detector models and derives the corresponding worldline amplitudes. Sec. V evaluates the excitation probability using a near-horizon stationary-phase analysis, isolating the universal thermal kernel and identifying the Proca-specific prefactors without invoking explicit formulas here. Sec. VI incorporates the cavity protocol and greybody transmission to promote probabilities to emission/absorption rates and to a master equation for the mode occupations. Sec. VII develops the entropy-flux analysis (HBAR) for massive spin-1 quanta and relates it to the black hole area change. Sec. VIII assembles consistency checks and theoretical subtleties, including control of the large-gap expansion, the role of the Proca constraint in place of gauge freedom, and the connection to Unruh/DeWitt detector logic. We close in Sec. IX with a brief summary and outlook.
@@ II Preliminaries and Conventions
We consider two-level atoms falling radially from rest at infinity into a nonrotating Schwarzschild black hole of mass $M$. The exterior quantum field is prepared in the Boulware vacuum. A mode-selecting cavity isolates ounterpropagating outgoing modes so that the atom-field interaction reduces to a single relevant frequency label $\nu$ for distant observers. Kinematics is encoded by the Schwarzschild line element and the tortoise coordinate $r_{*}$, and the atomic worldline is specified by $t(\tau)$ and $r(\tau)$ for proper time $\tau$. The near-horizon redshift makes the WKB phase of outgoing modes depend on the combination $t-r_{*}$, which will be the driver of the thermal detailed-balance factor later on.
@@ II.1 Geometry, trajectory, and coordinates
We work outside the Schwarzschild radius $r_{g}\equiv 2GM/c^{2}$ with line element [42, 19]
$ds^{2}=f(r)\,c^{2}dt^{2}-\frac{dr^{2}}{f(r)}-r^{2}(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}),$ (II.1)
with $f(r)=1-r_{g}/r$. For wave propagation, it is convenient to introduce the Regge-Wheeler tortoise coordinate [40, 26]
$r_{*}(r)=r+r_{g}\ln\!\left(\frac{r}{r_{g}}-1\right),$ (II.2)
so that outgoing high-frequency modes behave locally as $e^{-i\nu(t-r_{*}/c)}$ near the horizon (the precise field content will be specified later). In a small neighborhood of the horizon, one may use the Rindler approximation obtained by setting $r=r_{g}+z^{2}/(4r_{g})$ with $0<z\ll r_{g}$, which yields [9]
$ds^{2}\simeq\frac{z^{2}}{4r_{g}^{2}}\,c^{2}dt^{2}-dz^{2}-r_{g}^{2}d\Omega_{2}^{2},$ (II.3)
@@
Following Ref. [44], we adopt the dimensionless variables
$r$ $\rightarrow r_{g}\,r,$
$t$ $\rightarrow(r_{g}/c)\,t,$
$\omega$ $\rightarrow(c/r_{g})\,\omega,$
$\nu$ $\rightarrow(c/r_{g})\,\nu,$ (II.4)
@@
$\frac{dr}{d\tau}=-\frac{1}{\sqrt{r}},\qquad\frac{dt}{d\tau}=\frac{r}{r-1}\;,$ (II.5)
where $t$ is the Schwarzschild time and $\tau$ the atom’s proper time [both dimensionless in the convention of (II.4)]. Integrating (II.5), we obtain the parametric trajectory
$\tau(r)=-\frac{2}{3}\,r^{3/2}+{\rm const},$ (II.6)
$t(r)=-\frac{2}{3}\,r^{3/2}-2\sqrt{r}-\ln\!\left(\frac{\sqrt{r}-1}{\sqrt{r}+1}\right)+{\rm const},$ (II.7)
and the corresponding tortoise coordinate [40, 26]
$r_{*}(r)=r+\ln(r-1).$ (II.8)
Equations (II.5)–(II.8) completely specify $t(\tau)$, $r(\tau)$, and $r_{*}(\tau)$ for the ensuing stationary-phase analysis of the atom-field matrix elements. These expressions coincide with the kinematic relations used in Appendixes A and B of Scully et al. [44], to which we will refer for consistency of conventions.
We remark that the near-horizon Rindler form (II.3) rationalizes the universality of the phase $t-r_{*}$ and will underpin the Planckian factor arising later from the $\tau$-integral along the worldline. 2) The choice of additive constants in (II.6)–(II.7) fixes the zero of $t$ and $\tau$ and is immaterial for transition probabilities, which depend only on phase differences.
@@ II.2 Field state and cavity
We take the exterior quantum field to be in the Boulware vacuum, i.e. the vacuum defined by positive Killing frequency with respect to $\partial_{t}$ in the static exterior. Operationally, we place a one-dimensional, axially symmetric cavity along the radial direction to isolate a single, outgoing mode counterpropagating with the infalling atom cloud (the same role played by the cavity/mode selector in Scully et al [44]). The cavity suppresses contamination by incoming flux and renders the coarse-grained master-equation description well-posed; leakage can be neglected provided the photon loss rate is small compared to the atom-induced absorption rate.
@@
$a_{\nu,\lambda}\,|0_{B}\rangle=0,\qquad\nu>0,$ (II.11)
where $\lambda$ labels polarization (for the Proca field) or the single scalar polarization [10, 16]. This state choice is best viewed as an operational preparation of the cavity-filtered outgoing mode, not as a claim that the global exterior state of an evaporating black hole is Boulware. The purpose of the cavity is precisely to isolate a single outgoing channel with vanishing initial occupation, so that the atom’s counterrotating excitation probes the acceleration-radiation mechanism in its cleanest form. In this sense, the present calculation generalizes the original HBAR protocol to the Proca case.
By contrast, in the Unruh vacuum the outgoing sector carries a thermal occupation while the ingoing sector remains in vacuum; in the Hartle-Hawking vacuum both sectors are thermally populated. For those states, the near-horizon phase structure and the associated kernel derived in Secs. VVIII remain unchanged, but the cavity-mode master equation would acquire additional bath-dependent stimulated-emission and induced-absorption terms proportional to the corresponding occupation numbers. Thus, the Boulware calculation should be read as isolating the vacuum contribution to the selected outgoing channel, while alternative exterior states add state-dependent occupation effects on top of the same local kinematic kernel.
The near-horizon WKB form of each outgoing mode depends on the universal phase $t-r_{*}$; hence, when the atom falls through the cavity, the dominant contribution to the transition amplitude is governed by this phase, while the cavity merely enforces mode selection and provides the coarse graining needed for a laser-style master equation and the ensuing HBAR entropy analysis.